0.2[x-4]=x-5 over 5
Answer by stanbon(75887)

0.2[x-4]=x-5 over 5
------------
0.2(x-4) = (x-5)/5
(1/5)(x-4) = (1/5)(x-5)
x-4 = x-5
---------------
This is a contradiction.
There is no solution.
cheers,
Stan H.